# Classical $r$-matrix structure for elliptic Ruijsenaars chain and

### v1

@@ 1 Introduction and overview
In fact, the first elliptic model of type (1.7) was suggested by I. Krichever in [9] and was called the elliptic Toda lattice. The exact relation of (1.7) with this model requires several additional steps which we do not need in this paper. We will describe this relation in a separate publication.
@@
The first goal of this paper is to describe $r$-matrix structure of the Ruijsenaars chain (1.6)-(1.9). Namely, we compute the Poisson brackets $\{L^{a}_{1}(z),L^{b}_{2}(w)\}$ and represent them in the $r$-matrix form similarly to the original Ruijsenaars-Schneider model (1.5) corresponding to the case $n=1$. This allows us to calculate the brackets for the monodromy matrix $\{T_{1}(z),T_{2}(w)\}$, and the Poisson commutativity $\{{\rm tr}T^{k}(z),{\rm tr}T^{l}(w)\}=0$ follows from these results. The second goal is to show that the obtained $r$-matrix structure generalizes the Maillet non-ultralocal brackets (1.16) for the field analogue of the Calogero-Moser model (1.11)-(1.13). For this purpose we proceed to the level of field theory (1.17)-(1.18) and compute the continuous non-relativistic limit. We show that the brackets (1.16) are indeed reproduced in this limit.
@@ 2.2 Classical $r$-matrix
The classical $r$-matrix structure for the elliptic Ruijsenaars-Schneider model was proposed in [12]. The form of the Lax matrix in [12] differs from (2.1) by changing $b_{j}\rightarrow b_{i}$, that is these two Lax matrices are related by the gauge transformation with the diagonal gauge transformation matrix ${\rm diag}(b_{1},...,b_{N})$. For this reason the given below answer for the $r$-matrix structure slightly differs from the original one. For the reader’s convenience we recall the original $r$-matrix from [12] together with some standard notations in the Appendix B.
@@ 2.3 Non-relativistic limit to Calogero-Moser model
Summing up the above, we see that the term with $E_{1}(\eta)L^{\hbox{\tiny{RS}}}_{jj}(w)$ is cancelled:
$\begin{array}[]{c}\displaystyle{L^{\hbox{\tiny{RS}}}_{2}(w)s^{-}_{12}(z,w)=\sum\limits_{k=1}^{N}\sum\limits_{i\neq j}^{N}L^{\hbox{\tiny{RS}}}_{kj}(w)E_{1}(q_{ij}+\eta)E_{ii}\otimes E_{kj}+}\\ \displaystyle{+\sum\limits_{k\neq j}^{N}L^{\hbox{\tiny{RS}}}_{kj}(w)E_{1}(\eta)E_{jj}\otimes E_{kj}+\sum\limits_{j=1}^{N}L^{\hbox{\tiny{RS}}}_{jj}(w)E_{1}(w+\eta)E_{jj}\otimes E_{jj}}\end{array}$ (2.40)

### v2

@@ 1 Introduction and overview
In fact, the first elliptic model of type (1.7) was proposed by I. Krichever in [11] and called the elliptic Toda chain. Another model of this type is the Ruijsenaars-Toda chain introduced by V. Adler and Yu. Suris in [1]. The precise connection between (1.7) and these models requires several additional steps, which we will not need in this article. We will describe this relation in a separate publication.
The Ruijsenaars chain was derived in [22] in two ways. The first one is by gauge transformation of the XYZ chain, and the second way is by certain reduction from 2d Toda hierarchy. The original Ruijsenaars-Schneider model can also be obtained from 2d Toda lattice as well [12], and its hyperbolic version is obtained from the sine-Grodon model (two-periodic reduction from 2d Toda) [4, 16]. Below, we consider the non-relativistic limit, assuming the chain is relativistic. This terminology is taken from the original papers [17], where the term ”relativistic” arose due to the existence of the Poincaré algebra. In fact, we do not know if the chain model admits the Poincaré algebra. We use the term ”relativistic” to refer to (group-like) quadratic (in $L$) $r$-matrix structures, and the term ”non-relativistic” is used to refer to models described by linear (Lie algebra-like) $r$-matrix structures.
@@
The first goal of this paper is to describe $r$-matrix structure of the Ruijsenaars chain (1.6)-(1.9). Specifically, we compute the Poisson brackets $\{L^{a}_{1}(z),L^{b}_{2}(w)\}$ and represent them in $r$-matrix form similarly to the original Ruijsenaars-Schneider model (1.5) corresponding to the case $n=1$. This allows us to compute the brackets for the monodromy matrix $\{T_{1}(z),T_{2}(w)\}$, and these results imply the Poisson commutativity $\{{\rm tr}T^{k}(z),{\rm tr}T^{l}(w)\}=0$. The second goal is to show that the obtained $r$-matrix structure generalizes the Maillet non-ultralocal brackets (1.16) for the field analogue of the Calogero-Moser model (1.11)-(1.13). To do this, we move to the field theory level (1.17)-(1.18) and calculate the continuous non-relativistic limit. We show that the brackets (1.16) are indeed reproduced in this limit.
@@ 2.2 Classical $r$-matrix
The classical $r$-matrix structure for the elliptic Ruijsenaars-Schneider model was proposed in [15]. The form of the Lax matrix in [15] differs from (2.1) by substitution of $b_{j}\rightarrow b_{i}$, that is these two Lax matrices are related by the gauge transformation with the diagonal gauge transformation matrix ${\rm diag}(b_{1},...,b_{N})$. For this reason, the answer given below for the $r$-matrix structure differs slightly from the original. For the reader’s convenience, we recall the original r-matrix from [15] along with some standard notation in Appendix B.
@@ 2.3 Non-relativistic limit to Calogero-Moser model
To summarize the above, we see that the term with $E_{1}(\eta)L^{\hbox{\tiny{RS}}}_{jj}(w)$ cancels:
$\begin{array}[]{c}\displaystyle{L^{\hbox{\tiny{RS}}}_{2}(w)s^{-}_{12}(z,w)=\sum\limits_{k=1}^{N}\sum\limits_{i\neq j}^{N}L^{\hbox{\tiny{RS}}}_{kj}(w)E_{1}(q_{ij}+\eta)E_{ii}\otimes E_{kj}+}\\ \displaystyle{+\sum\limits_{k\neq j}^{N}L^{\hbox{\tiny{RS}}}_{kj}(w)E_{1}(\eta)E_{jj}\otimes E_{kj}+\sum\limits_{j=1}^{N}L^{\hbox{\tiny{RS}}}_{jj}(w)E_{1}(w+\eta)E_{jj}\otimes E_{jj}\,.}\end{array}$ (2.40)